The exercise involving data in this and subsequent sections were designed to be solved using Excel.

The following estimated regression equation is based on 10 observations was presented.

y=29.1270+.5906x1+.498x2

Here SST=6566.250 SSR=6046.625 Sb1=.0821 and Sb2=.0672

1. Compute MSR and MSE to 3 decimal places.

2. Compute the F statistic to 2 decimal places.

2a. What is the p-value?

3. Compute the t test statistic for the significance of b1(2 decimals)

3a. What is the p-value?

4. Compute the t test statistic for the significance of b2(2 decimals)

4a. What is the p-value?
Answer #1

1)

p=number of independent variables= 2
MSR =SSR/p = 3023.313
MSE=(SSE/(n-p-1))= 74.232

2)

F test statistic =MSR/MSE= 40.73
p value = 0.0001             or p value <0.01

3)

t test statistic =coefficient/standard error = 7.19
p value = 0.0002             or p value <0.01

4)

t test statistic =coefficient/standard error = 7.41
p value = 0.0001             or p value <0.01